Say true or false, with a short proof.


All the cube roots of I in C are z₁ = cos(π/2)+ i sin(π/2) , z₂ = cos (π/6)+ isin(π/6) and z₃ = cos (5π/6)+i sin (5π/6)


1
Expert's answer
2021-08-29T16:40:55-0400

The polar form of "i" is "1(\cos(\dfrac{\pi}{2})+i\sin(\dfrac{\pi}{2}))."

According to the De Moivre's Formula, all cubic roots of a complex number "1(\cos(\dfrac{\pi}{2})+i\sin(\dfrac{\pi}{2}))" are given by "1^{1/3}(\cos(\dfrac{\pi/2+2\pi k}{3})+i\sin(\dfrac{\pi/2+2\pi k}{3})), k=0,1,2."


"k=0, \sqrt[3]{1}(\cos(\dfrac{\pi/2+2\pi (0)}{3})+i\sin(\dfrac{\pi/2+2\pi (0)}{3}))"

"=\cos(\dfrac{\pi}{6})+i\sin(\dfrac{\pi}{6})"



"k=1, \sqrt[3]{1}(\cos(\dfrac{\pi/2+2\pi (1)}{3})+i\sin(\dfrac{\pi/2+2\pi (1)}{3}))"

"=\cos(\dfrac{5\pi}{6})+i\sin(\dfrac{5\pi}{6})"




"k=2, \sqrt[3]{1}(\cos(\dfrac{\pi/2+2\pi (2)}{3})+i\sin(\dfrac{\pi/2+2\pi (2)}{3}))"

"=\cos(\dfrac{3\pi}{2})+i\sin(\dfrac{3\pi}{2})"


Therefore the statement is False.